A 23-year, semiannual coupon bond sells for $981.73. The bond has a par value of $1,000 and a yield to maturity of 6.81 perceA 23-year, semiannual coupon bond sells for $981.73. The bond has a par value of $1,000 and a yield to maturity of 6.81 percent. What is the bond's coupon rate?
Answer #1

price of coupon = Coupon payment per period * [1-(1+i)^-n]/i + par value/(1+i)^n

i = interest rate per period

n = number of periods

=>

981.73 = C/2 * [1-(1+0.0681/2)^-46]/(0.0681/2) + 1000/(1+0.0681/2)^46

=>

coupon payment = 66.52

hence

coupon rate = 66.52/1000 = 6.65%